# Congestion or No Congestion: Packet Loss Identification and Prediction Using Machine Learning

### v2

@@ I Introduction
Subsequent sections of this paper are organized as follows: Section 2 provides a brief overview of the literature on loss prediction and congestion management. In Section 3, we discuss the problem statement in detail. We explain the methodology used in our study in section 4. Section 5 presents and discusses the results. Finally, we conclude the paper in Section 6.
@@ V-A Machine learning models
Random Forest Classifier qDrop 0.94 0.97 1233 1158 0.88 1158 53 22
wDrop 0.62 0.67 251 212 0 156 95
K-Neighbor Classifier qDrop 0.94 0.97 1233 1158 0.89 1158 75 0
wDrop 0.75 0.70 251 292 0 189 62
Gradient Boosting Classifier qDrop 0.72 0.83 1233 883 0.84 808 25 1
wDrop 0.63 0.68 251 211 3 678 50
Logistic Reg Classifier qDrop 0.95 0.97 1233 1184 0.88 809 15 10
wDrop 0.62 0.68 251 205 39 607 85
Decision Tree Classifier qDrop 0.72 0.83 1233 883 0.77 883 269 81
wDrop 0.67 0.47 251 457 0 167 84
@@ VI Results
The Macro Average F1 scores, which average the F1-scores across both classes, indicate that RF and KNN classifiers are the most balanced across packet types, with the highest scores of $0.88$ and $0.89$, respectively. The Confusion Matrix highlights the specific areas of misclassification for each model, with RF and KNN showing fewer misclassifications compared to the other models. Overall, the RF and KNN models emerged as the most effective classifiers for this task, with better overall performance in classifying packet loss types in the network traffic data.

### v3

@@ I Introduction
Subsequent sections of this paper are organized as follows: Section 2 provides a brief overview of the literature on loss prediction and congestion management. In Section 3, we discuss the problem statement in detail. We explain the dataset in section 4. Section 5 presents the methodology used in this study. We discuss the results in section 6. Finally, we conclude the paper in Section 7.
@@ V-A Machine learning models
Random Forest Classifier qDrop 0.94 0.97 1233 1158 0.88 1158 0 0
wDrop 0.62 0.67 251 212 53 156 3
K-Neighbor Classifier qDrop 0.94 0.97 1233 1158 0.89 1158 0 0
wDrop 0.75 0.70 251 292 75 189 28
Gradient Boosting Classifier qDrop 0.72 0.83 1233 883 0.84 883 0 0
wDrop 0.63 0.68 251 211 48 158 5
Logistic Reg Classifier qDrop 0.95 0.97 1233 1184 0.88 1174 10 0
wDrop 0.62 0.68 251 205 35 155 15
Decision Tree Classifier qDrop 0.72 0.83 1233 883 0.77 883 0 0
wDrop 0.67 0.47 251 472 273 173 26
@@ VI Results
The Macro Average F1 scores, which average the F1-scores across both classes, indicate that RF and KNN classifiers are the most balanced across packet types, with the highest scores of $0.88$ and $0.89$, respectively. The Confusion Matrix highlights the specific areas of misclassification for each model, with RF, LR, and KNN showing fewer misclassifications compared to the other models. Overall, the RF and KNN models emerged as the most effective classifiers for this task, with better overall performance in classifying packet loss types in the network traffic data.